OpenStudy (anonymous):

A small puck with a mass of 0.10 kilograms is sliding in the +x-direction on an air table with an initial speed of 8.0 meters/second. The puck collides with a larger puck that is initially at rest. The larger puck has a mass of 0.40 kilograms. After the collision, the smaller puck has a speed of 3.4 meters/second at an angle of +60° from its original path. The angle made by the larger puck is -35°. What is the x-momentum after the collision if the velocity of the larger puck after the collision is 1.5 meters/second?

OpenStudy (anonymous):

Make a sketch of the situation first to help you understand the problem. So the momentum is speed*mass=0,1*8=0.8 (for the small puck before collision) After collision the momentum of the small puck is 0.1*3.4=0.34 The rest of the momentum goes towards the bigger one 0.4*1.5=0.6 This is simply stupid as it gained momentum by colliding. That is just impossible.

OpenStudy (anonymous):

For the x momentum you just need a bit of trigonometry

OpenStudy (anonymous):

So I'll be working with .8, .34, and .6? Is there a formula I can use ?